# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

### v6

@@ 6 Performance
The experiments were conducted using the settings outlined in Tab. 8 of Appendix E. Time costs were recorded on an NVIDIA A100 40GB GPU.
@@
Figure 4: Perturbations of adversarial attacks and the proposed preemptive defenses. Visual comparison between the original image and the RE demonstrates that Fast Preemption perturbations are imperceptible. Additionally, comparison of perturbations between AutoAttack and Fast Preemption reveals that preemptive defense is more likely to succeed when perturbations are similar.
@@ 6.1 Robustness against Adversarial Attacks
We evaluated transferable defense performance under the condition where the defense lacks knowledge of both the attacks and victim models [6, 29]. Table 1 and Tab. 2 compares Fast Preemption with benchmark defenses against four attacks. Fast Preemption outperformed the benchmarks in both clean and robust accuracies while maintaining competitive time costs. For instance, compared to the least time-costly method, WideResNet-94-16 [6] on CIFAR-10 under Standard $L_{\infty}$ AutoAttack, Fast Preemption achieved 1.6% and 11.9% higher clean and robust accuracies, respectively, with only a marginal increase of 0.04 seconds per generation. Compared to the SOTA preemptive defense Bi-Level [32], which had the second-best accuracy, Fast Preemption improved clean and robust accuracies by 2.5% and 5.1%, respectively, while being 700 times faster. Notably, preemptive defenses can be combined with training- and test-time defenses. For example, when paired with a test-time defense, Fast Preemption’s performance was significantly enhanced, as shown in the final rows of CIFAR-10 and ImageNet in Tab. 1.
Figure 4 provides a qualitative analysis, comparing the original images with their corresponding REs. The perturbations introduced by Fast Preemption are barely perceptible. The average SSIM and LPIPS scores for 2,000 examples from CIFAR-10 and ImageNet are 0.97 and 0.01, respectively, both surpassing those of standard AutoAttack and the SOTA preemptive defense Bi-Level [32] (more details in Appendix A). Additionally, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations closely resemble those of the attack. The perturbation illustration $\delta_{Gray}$ was generated by normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$, converting $\delta_{RGB}$ to grayscale using the NTSC formula [30].
@@ 6.2 Resistance against Reversion Attacks
We conducted reversion attacks to assess the reversibility of preemptive defenses. In addition to the proposed Preemptive Reversion attack, we employed JPEG compression and adversarial purification, which do not require knowledge of the defense. Preemptive Reversion was tested in both white-box and black-box scenarios (see Sec. 4.2). The results, summarized in Tab. 3, reveal that only Preemptive Reversion, indicated in red, succeeded in reversing the preemptive defense, bringing the accuracy closer to the original. Other reversion attacks failed, further degrading clean accuracy due to distortion. Notably, Preemptive Reversion is only effective in the white-box setting, as Fast Preemption remains irreversible without full knowledge of the defense.
Figure 5 corroborates our findings, showing that the white-box reversion successfully eliminates most protective perturbations, whereas the black-box reversion fails. However, due to the randomization inherent in Fast Preemption (i.e., randomly initialized FGSM), the white-box reversion cannot entirely remove all protective perturbations. The fourth row in Tab. 3, which was obtained without our evaluation protocol, further highlights the comprehensiveness of our evaluation method. The other results demonstrate that Preemptive Reversion in the white-box setting is the only reversion attack that improves clean accuracy after reversion. The lower numbers in the fourth row suggest that the accuracy decrease observed with Preemptive Reversion in the white-box setting resulted from successful reversion, while other accuracy declines were due to distortion. Consequently, a decrease in accuracy, when solely considered from the fourth row, does not necessarily indicate the success of a reversion attack.
@@ 7 Ablation Studies
We conducted an ablation study on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset and AutoAttack (Standard) [11] ($L_{\infty}$-norm $\epsilon=8/255$) against a black-box target model, RaWideResNet-70-16 [35]. As shown in Fig. 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [32], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to 0.04 seconds. These studies were conducted to identify optimal configurations, with further ablation results provided in Appendix B.
@@ 8 Conclusion
We propose Fast Preemption, an efficient and transferable preemptive adversarial defense that introduces a new layer in the defense hierarchy: Proactive Defense (ours) → Training-Time Defense → Test-Time Defense. Unlike conventional training- and test-time defenses, Fast Preemption is plug-and-play, transferable, and requires no prior knowledge, enabling users to safeguard media independently without relying on third-party interventions. Our approach also addresses the efficiency limitations of existing proactive defenses, achieving processing speeds 700 times faster than prior methods, an essential improvement for real-world applications, where significant delays (e.g., 28 seconds per image [32]) are impractical. Moreover, Fast Preemption maintains or improves clean accuracy while achieving SOTA robust accuracy, overcoming the common trade-off where defenses degrade clean accuracy.
Despite its advantages, our defense exhibits limitations against the proposed reversion attack in the white-box setting. While allowing users to customize the backbone model as a user-side key mitigates reversion attacks, further work is needed to simplify user requirements. Future research should also investigate the relationship between backbone models and transferability, as discussed in Sec. B.2.
@@
This supplementary material provides a comprehensive analysis of visual quality (Appendix A) and presents an extensive ablation study (Appendix B), investigating critical components such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples. Additionally, it includes evaluation results against a broader set of models (Appendix C), elaborates on the algorithms underlying the proposed adversarial defense and adaptive reversion attack (Appendix D), and details the experimental configurations for both attack and defense scenarios (Appendix E).
Appendix A Visual Quality
Figure 7: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA proactive defense [32]. All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images. SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
A.1 Structural Similarity Index (SSIM)
where:
  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
  • $\sigma_{xy}$: Covariance between $x$ and $y$.
    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
The SSIM value ranges from -1 to 1, where:
  • 1: Perfect similarity.
  • 0: No similarity.
  • Negative values indicate dissimilarity.
A.2 Learned Perceptual Image Patch Similarity (LPIPS)
$LPIPS(x,y)=$ (17)
where:
  • $l$: Layer index in the deep network.
  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
Appendix B Ablation Studies
Table 4: Performance with different classifiers.
       Ablation        Time Cost(s)        Transferable Defense
       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
       Ground Truths        100        28        97.2        86.9        83.1
       ResNeXt        95.9        28        95.4        85.7        82.7
       Bi-Level        89.1        29        91.9        83.9        81.0
B.1 Classifier
In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Tab. 4, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
Table 5: Performance of different backbone models.
Backbone Robustness Clean (%) 89.1 93.9 88.1
PGD (%) 43.9 77.1 61.8
AutoAttack (%) 42.4 74.0 58.6
Time Cost (s) for Defense 28 69 26
White-Box Defense Clean (%) 95.5 95.8 93.8
PGD (%) 84.9 91.6 85.9
AutoAttack (%) 81.8 90.2 83.3
PGD (%) 85.7 89.7 87.6
AutoAttack (%) 82.7 88.0 85.9
Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
PGD (%) 72.7 78.6 79.6
AutoAttack (%) 70.8 75.7 77.3
B.2 Backbone Model
B.3 Forward-Backward Cascade Learning
Table 6: Performance for different $N$ and optimization algorithms.
       Ablation        Time Cost (s)        Transferable Defense
       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
       PGD        1        0.7        95.3        87.5        85.8
       FGSM        1        0.04        95.4        86.8        85.4
       PGD        20        0.8        95.2        87.7        85.9
       FGSM        20        0.04        95.5        87.8        85.9
B.4 Optimization Algorithm
Table 7: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. All proactive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
Model Metric No Proactive Defense Bi-Level [32] Fast Preemption (ours)
Time Cost (s) N/A 28 0.35 0.04
RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
Robust 71.4 81.3 86.8 85.9
WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
Robust 71.2 80.9 86.4 86.0
WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
Robust 62.1 76.2 78.7 80.8
WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
Robust 64.8 77.7 81.3 83.2
PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
Robust 42.4 54.7 55.4 61.2
ResNest152 [41] Clean 86.9 87.6 91.2 92.4
Robust 62.4 73.2 75.8 78.2
ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
Robust 52.0 68.1 66.7 72.0
ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
Robust 51.4 67.5 67.8 72.8
XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
Robust 57.9 71.3 74.6 76.0
Average Clean 89.9 89.8 93.4 94.1
Robust 59.5 72.3 74.8 77.3
  • Bold indicates the best performance.
Appendix C Evaluation across More Models
As shown in Tab. 7, our method demonstrates significant performance gains compared to no proactive defense and the SOTA proactive defense [32]. Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively. Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
Appendix D Detailed Algorithms
@@
Appendix E Experiment Settings
@@
Appendix F Related Works
Existing adversarial defenses are divided into three categories: training-time, test-time, and preemptive.
Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels. The earliest mechanisms, e.g., adversarial training [18], incorporated various types of AEs along with their ground-truth labels in the training dataset. Zhang et al. [54] proposed TRADES, which introduces a theoretically grounded approach to balance robustness and accuracy in adversarial training by incorporating a custom loss function that controls the trade-off between clean accuracy and adversarial robustness. To improve the efficiency of training on large-scale datasets, Shafahi et al. [42] devised accelerated adversarial training in which backward-pass computations are reused. This entails curating a large number of specific real-world samples; Sehwag et al. [41] circumvented this problem by using additional data from proxy distributions learned by advanced generative models for adversarial training. Gowal et al. [19] demonstrated that generative models trained solely on the original dataset can be leveraged to artificially increase the training dataset’s size and improve adversarial robustness against $L_{p}$-norm perturbations. Debenedetti et al. [13] proposed using a custom recipe of vision transformers for adversarial training in which weaker data augmentation is used with warm-up in the augmentation intensity. Inspired by denoising diffusion models [23], Wang et al. [50] conducted adversarial training using data generated by an elucidating diffusion model [24]. Dong et al. [14] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration. Peng et al. [35] devised a robust model which follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions. Bartoldson et al. [6] developed the current state-of-the-art (SOTA) robust model by scaling laws for adversarial training to assess robustness in image classifiers, revealing that even with significant computational resources, robustness asymptotes below human-level performance due to flaws in attack formulations generating invalid data, suggesting future efforts should address both computational efficiency and image validity.
Training-time defenses have two shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al. [14], which was designed to circumvent this problem.
Test-time defenses are divided to detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs. Among recent studies, Cohen et al. [9] devised a k-nearest neighbor model for ranking the most supportive training samples for any given validation example in training a detector. Zuo and Zeng [56] proposed using a binary classifier, erase-and-restore, that uses the intriguing sensitivity of $L_{2}$-norm attacks by randomly erasing certain pixels in an $L_{2}$-norm AE and then restoring them with an inpainting technique. Liang et al. [27] detected AEs by comparing the classification results of a given sample and a denoised version that was generated through adaptive noise reduction combining scalar quantization and a smoothing spatial filter. Most recently, Hickling et al. [22] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks [7]. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification. For instance, Wang et al. [48] distorted images by randomly nullifying pixels. Guo et al. [20] preprocessed inputs with a nondifferentiable transformation. Raff et al. [37] combined a large number of individually weak defenses into a single randomized transformation to build a strong defense. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious. Ren et al. [38] trained a generative model using autoencoder technology [25]. They claimed the existence of an immune space, in which perturbations have fewer adverse effects, and they inactivated adversarial perturbations by restricting them to that subspace. Nie et al. [34] devised the DiffPure benchmark, which uses denoising diffusion models [23]. DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [46].
Lastly, Alfarra et al. [3] introduced a test-time defense mechanism incorporating preemptive defense as an anti-adversary layer guided by the output of a standard classifier. However, this defense can be circumvented if the classifier is compromised, a limitation explicitly acknowledged in their paper. This significant vulnerability undermines the practicality of using preemptive defense techniques in test-time scenarios. While they characterize this scenario as a “least realistic setting,” we contend that it aligns with the standard white-box attack paradigm, where adversaries have full access to the model. Under even more challenging assumptions, adversaries might also be aware of the defense mechanism itself (white-box reversion). Furthermore, the approach of Alfarra et al. [3] depends exclusively on forward propagation-based learning, which increases the risk of overfitting.
Preemptive defenses aim to generate REs that replace original images, which are then discarded. These REs are more challenging to manipulate and thus resist adversarial attacks. Salman et al. [40] was the first to devise a preemptive defense strategy utilizing REs, which they referred to as “unadversarial” examples. They were learned using projected gradient descent (PGD) [31] and improved both in-distribution performance and robustness against unforeseen perturbations. Moon et al. [32] devised a Bi-Level optimization algorithm (abbreviated as Bi-Level in this paper) that finds points in the vicinity of natural images that are robust against adversarial perturbations; it achieved SOTA robust accuracy. However, its time cost is huge, and its transferability is low. Frosio and Kautz [17] introduced the $A^{5}$ (“Adversarial Augmentation Against Adversarial Attacks”) framework that uses an individual classifier and a robustifier (a generator), which are co-trained. It is much more efficient than the Bi-Level optimization algorithm, but it is not as robust.

### v7

@@ 6 Performance
The experiments were conducted using the settings outlined in Tab. 4 of Appendix B. Time costs were recorded on an NVIDIA A100 40GB GPU.
@@ 6.1 Robustness against Adversarial Attacks
We evaluated transferable defense performance under the condition, where the defense lacks knowledge of both the attacks and victim models [6, 26]. Tab. 1 and Tab. 2 compares Fast Preemption with benchmark defenses against four attacks. Fast Preemption outperformed the benchmarks in both clean and robust accuracies while maintaining competitive time costs. For instance, compared to the least time-costly method, WideResNet-94-16 [6] on CIFAR-10 under Standard $L_{\infty}$ AutoAttack, Fast Preemption achieved 1.6% and 11.9% higher clean and robust accuracies, respectively, with only a marginal increase of 0.04 seconds per generation. Compared to the SOTA preemptive defense Bi-Level [29], which had the second-best accuracy, Fast Preemption improved clean and robust accuracies by 2.5% and 5.1%, respectively, while being 700 times faster. Notably, preemptive defenses can be combined with training- and test-time defenses. For example, when paired with a test-time defense, Fast Preemption’s performance was significantly enhanced, as shown in the final rows of CIFAR-10 and ImageNet in Tab. 1.
Figure 4: Perturbations of adversarial attacks and preemptive defenses. Visual comparison between the original image and the robust example (RE) shows that Fast Preemption perturbations are imperceptible. Furthermore, comparison of perturbations between AutoAttack and Fast Preemption indicates that the preemptive defense is more likely to succeed when perturbations are similar.
Fig. 4 provides a qualitative analysis, comparing the original images with their corresponding REs. The perturbations introduced by Fast Preemption are barely perceptible. The average SSIM and LPIPS scores for 2,000 examples from CIFAR-10 and ImageNet are 0.97 and 0.01, respectively, both surpassing those of AutoAttack and the SOTA preemptive defense Bi-Level [29] (more details in Appendix D). Additionally, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations closely resemble those of the attack. The perturbation illustration $\delta_{Gray}$ was generated by normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$, converting $\delta_{RGB}$ to grayscale using the NTSC formula [27].
@@ 6.2 Resistance against Reversion Attacks
We conducted reversion attacks to assess the reversibility of preemptive defenses. In addition to the proposed Preemptive Reversion attack, we employed JPEG compression and adversarial purification, which do not require knowledge of the defense. Preemptive Reversion was tested in both white-box and black-box scenarios (see Sec. 4.2). The results, summarized in Tab. 3, reveal that only Preemptive Reversion, indicated in red, succeeded in reversing the preemptive defense, bringing the accuracy closer to the original. Other reversion attacks failed, further degrading clean accuracy due to distortion.
Notably, Preemptive Reversion is effective only in the white-box setting, which is less realistic for our preemptive defense, as our defense allows confidential user configurations (e.g., selection of the backbone model). In the black-box setting, where the backbone models are unknown, all adaptive attacks fail to bypass the preemptive defense, further demonstrating its reliability. In contrast, other preemptive defenses, such as Bi-Level [29], rely on paired protection-inference models, meaning that compromising the target model is sufficient to reverse user protections.
Fig. 5 corroborates our findings, showing that the white-box reversion successfully eliminates most protective perturbations, whereas the black-box reversion fails. However, due to the randomization inherent in Fast Preemption (i.e., randomly initialized FGSM), the white-box reversion cannot entirely remove all protective perturbations. The fourth row in Tab. 3, which was obtained without our evaluation protocol, further highlights the comprehensiveness of our evaluation method. The other results demonstrate that Preemptive Reversion in the white-box setting is the only reversion attack that improves clean accuracy after reversion. The lower numbers in the fourth row suggest that the accuracy decrease observed with Preemptive Reversion in the white-box setting resulted from successful reversion, while other accuracy declines were due to distortion. Consequently, a decrease in accuracy, when solely considered from the fourth row, does not necessarily indicate the success of a reversion attack.
@@ 7 Ablation Studies
We conducted an ablation study on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset and AutoAttack (Standard) [11] ($L_{\infty}$-norm $\epsilon=8/255$) against a black-box target model, RaWideResNet-70-16 [32]. As shown in Fig. 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [29], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to 0.04 seconds. These studies were conducted to identify optimal configurations, with further ablation results provided in Appendix E.
@@ 8 Conclusion
We propose Fast Preemption, an efficient, transferable, plug-and-play preemptive adversarial defense that introduces a new layer in the defense hierarchy: Preemptive Defense (ours) → Training-Time Defense → Test-Time Defense. Fast Preemption minimizes the need for prior knowledge, allowing users to independently safeguard media without relying on third-party interventions. It achieves SOTA clean and robust accuracy among training-time, test-time, and preemptive defenses while addressing the efficiency limitations of existing preemptive approaches, achieving processing speeds up to 700 times faster.
Additionally, we propose Preemptive Reversion, a more robust reversion attack that demonstrates our preemptive defense’s resilience against reversion by incorporating a strategy that maintains the confidentiality of user configurations.
However, our defense exhibits limitations against the proposed reversion attack. While allowing users to customize defense configurations mitigates reversion attacks, further work aims to simplify user requirements. Future research should also investigate the relationship between backbone models and transferability.
@@
This supplementary material provides a detailed exposition of the algorithms underlying the proposed adversarial defense and adaptive reversion attack (Appendix A), outlines the experimental configurations for both attack and defense scenarios (Appendix B), presents evaluation results against a broader range of models (Appendix C), and offers a comprehensive analysis of visual quality (Appendix D). Additionally, it includes an extensive ablation study (Appendix E), examining key components such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the effectiveness of averaging multiple Fast Gradient Sign Method (FGSM) examples.
Appendix A Detailed Algorithms
@@
Appendix B Experiment Settings
@@
Appendix C Evaluation across More Models
As shown in Tab. 6, our method demonstrates significant performance gains compared to no preemptive defense and the SOTA preemptive defense [29]. Specifically, we achieve average improvements of 4.2% and 4.3% in clean accuracy, 17.8% and 5% in robust accuracy, and a 700-fold increase in speed, respectively. Furthermore, compared using the SOTA adversarial training model [6] as the backbone model, our chosen [19] yields an average increase of 0.7% in clean accuracy, 2.5% in robust accuracy, and a tenfold improvement in processing speed.
Table 6: Transferable clean and robust accuracy (%) across various victim models against AutoAttack (Standard) [11] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. All preemptive defenses were conducted using classifier and backbone models distinct from the victim models, even when the architectures were similar.
Model Metric No Preemptive Defense Bi-Level [29] Fast Preemption (ours)
Time Cost (s) N/A 28 0.35 0.04
RaWideResNet-70-16 [32] Clean 93.5 92.1 95.5 95.4
Robust 71.4 81.3 86.8 85.9
WideResNet-70-16 [47] Clean 93.3 91.5 95.5 95.3
Robust 71.2 80.9 86.4 86.0
WideResNet-34-10 [33] Clean 91.5 90.2 95.1 95.1
Robust 62.1 76.2 78.7 80.8
WideResNet-28-10 [49] Clean 93.4 91.5 95.4 95.6
Robust 64.8 77.7 81.3 83.2
PreActResNet-18 [48] Clean 84.2 87.3 90.1 91.6
Robust 42.4 54.7 55.4 61.2
ResNest152 [38] Clean 86.9 87.6 91.2 92.4
Robust 62.4 73.2 75.8 78.2
ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
Robust 52.0 68.1 66.7 72.0
ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
Robust 51.4 67.5 67.8 72.8
XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
Robust 57.9 71.3 74.6 76.0
Average Clean 89.9 89.8 93.4 94.1
Robust 59.5 72.3 74.8 77.3
  • Bold indicates the best performance.
Appendix D Visual Quality
Figure 7: Visual quality of robust examples (REs) generated by our Fast Preemptive defense, compared with adversarial examples (AEs) from AutoAttack and REs from the SOTA preemptive defense [29]. All samples adhere to the same $L_{\infty}$-norm constraint ($\epsilon=8/255$) relative to the original images. SSIM and LPIPS scores demonstrate that the visual quality of our REs is sufficiently high to serve as replacements for the original images, and is comparable to (and often slightly better than) those of the adversarial attack and existing defense methods.
D.1 Structural Similarity Index (SSIM)
where:
  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
  • $\sigma_{xy}$: Covariance between $x$ and $y$.
    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
The SSIM value ranges from -1 to 1, where:
  • 1: Perfect similarity.
  • 0: No similarity.
  • Negative values indicate dissimilarity.
D.2 Learned Perceptual Image Patch Similarity (LPIPS)
$LPIPS(x,y)=$ (17)
where:
  • $l$: Layer index in the deep network.
  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
Appendix E Ablation Studies
Table 7: Performance with different classifiers.
       Ablation        Time Cost(s)        Transferable Defense
       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
       Ground Truths        100        28        97.2        86.9        83.1
       ResNeXt        95.9        28        95.4        85.7        82.7
       Bi-Level        89.1        29        91.9        83.9        81.0
E.1 Classifier
In all preemptive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Tab. 7, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
Table 8: Performance of different backbone models.
Backbone Robustness Clean (%) 89.1 93.9 88.1
PGD (%) 43.9 77.1 61.8
AutoAttack (%) 42.4 74.0 58.6
Time Cost (s) for Defense 28 69 26
White-Box Defense Clean (%) 95.5 95.8 93.8
PGD (%) 84.9 91.6 85.9
AutoAttack (%) 81.8 90.2 83.3
PGD (%) 85.7 89.7 87.6
AutoAttack (%) 82.7 88.0 85.9
Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
PGD (%) 72.7 78.6 79.6
AutoAttack (%) 70.8 75.7 77.3
E.2 Backbone Model
E.3 Forward-Backward Cascade Learning
Table 9: Performance for different $N$ and optimization algorithms.
       Ablation        Time Cost (s)        Transferable Defense
       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
       PGD        1        0.7        95.3        87.5        85.8
       FGSM        1        0.04        95.4        86.8        85.4
       PGD        20        0.8        95.2        87.7        85.9
       FGSM        20        0.04        95.5        87.8        85.9
E.4 Optimization Algorithm